OpenStudy (anonymous):

If x^2+y^2=25 then find an expression for y'' Steps please, so i can compare to mine to see if i did this correct? I have a quiz tmr. I got (25/y^3) but i have a feeling i messed up on a sign somewhere.

OpenStudy (anonymous):

i just need the steps for the second part where we find y'' and plug in the -25 at the end

OpenStudy (anonymous):

\[2xx'+2yy'=0\] \[y'=-\frac{xx'}{y}\] and then repeat

OpenStudy (anonymous):

im asking for specifics :3 lol i know the deriv. rules

zepdrix (zepdrix):

Derivative with respect to x?

OpenStudy (anonymous):

no to y

OpenStudy (anonymous):

\[25-x ^{2} = y ^{2}\] \[y = \sqrt{25 - x ^{2}}\]

zepdrix (zepdrix):

no you're looking for derivative with respect to y. it says y'' at the top. maybe you misunderstood what i was asking :o

zepdrix (zepdrix):

respect blahhhh to x.. typo -_-

OpenStudy (anonymous):

the second derivative

zepdrix (zepdrix):

 hmm this is what i came up with